OpenStudy (anonymous):

Simplify this problem:

OpenStudy (anonymous):

\[2/\sqrt{5}\]

OpenStudy (lifeisadangerousgame):

This actually belongs in the mathematics section ;)

OpenStudy (kirbykirby):

I'll help anyway. I believe they want you to rationalize the fraction. So multiply the the square root to the numerator and denominator (just basically multiplying by 1). \[\frac{2}{\sqrt5} = \frac{2}{\sqrt5}*\frac{\sqrt5}{\sqrt5} = \frac{2\sqrt5}{5}\] So you end up just multiplying the numerator part and denominator part just like regular fractions. The end result is to get a fraction that doesn't have a square root.